OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5 A. 0 B. 1 C. 2 D. infinite

OpenStudy (texaschic101):

y = -1/2x + 5 (change to standard form) 1/2x + y = 5 -- multiply by 2 x + 2y = 10 -2x + 4y = -7 should be 1 solution

OpenStudy (anonymous):

@texaschic101 Thank you so much.

OpenStudy (texaschic101):

sure thing :)